Syazana inherits RM25 000 in April 2012. She invests the money at an interest rate of

2.7% per year. What is the value of her investment in April 2025?


1
Expert's answer
2021-09-27T15:55:37-0400

Compound interest formula


"A=P(1+\dfrac{r}{n})^{nt}"

"P=RM25000"

"n=1"

"r=0.027"

"t=2025-2012=13"


"A=25000(1+\dfrac{0.027}{1})^{1(13)}=35347.26"

Her investment in April 2025 will be "RM35347.26"